OpenStudy (anonymous):

what are the steps to : 0=m/4-1

OpenStudy (anonymous):

tell me if I'm understanding this correctly \[0 = \frac{ m }{ 4-1 }\] it that right?

OpenStudy (anonymous):

or is it: \[0 = \frac{ m }{ 4 } - 1\]

OpenStudy (anonymous):

no, its: 0=\[\frac{ m }{ 4 }\]-1

OpenStudy (anonymous):

so like the second one i wrote there?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok, so we start by adding 1 to both sides of the equation, like this: \[0+1 = \frac{ m }{ 4 } - 1 + 1\] you can see that the positive and negative 1 on the right hand side of the equation cancel each other out, and 0+1 just equal 1, so we re-write as: \[1 = \frac{ m }{ 4 }\]

OpenStudy (anonymous):

following this ok so far, lemme know if theres anything u dont get

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so then you can multiply both sides by 4 to get: \[1 \times 4 = \frac{ m \times 4 }{ 4 }\] here, the 4's on the right hand side of the equation cancel each other out, and 1 times 4 is just 4 so: \[4 = m\]